# SANSA: Unleashing the Hidden Semantics in SAM2 for Few-Shot Segmentation

### v1

@@ 3.2 Feature Adaptation
We hypothesize that SAM2 features already encode semantic concepts, albeit entangled with signals specific to tracking, such as instance-level details and spatial biases. If such structure exists, then it should be learnable from a set of base classes by training few parameters [1, 27, 51]. Thus, we opt for simple AdaptFormer [11] blocks, although our analysis is not tied to the adaptation method, as we will show in the Experimental Section. We integrate AdaptFormer blocks within the last two layers of the Image Encoder, as these encode higher-level semantic representations. Given down- and up-projection matrices $\mathbf{W}_{down}\in\mathcal{R}^{d,\tilde{d}},\mathbf{W}_{up}\in\mathcal{R}^{\tilde{d},d}$, an AdaptFormer block operates token-wise:
@@ 3.3 Training objective
Following standard practice in FSS, we adopt an episodic training paradigm [60, 53, 62, 31]. We have access to a set of training episodes, each one containing annotated instances from a single class. We denote these episodes as $\{x^{i},y^{i}\}$, where $y^{i}\in[0,1]^{H\times W}$ is the mask that segments these instances.
@@
where $\{x_{r},a_{r}\}$ is the annotated reference image and $\{x^{j}_{t}\}_{j=1}^{J}$, are the $J$ unlabeled target images. We feed $\mathcal{M}_{train}$ to our model, which propagates the provided concept across target frames to predict the masklet $\{\hat{y}^{j}_{t}\}_{j=1}^{J}$. Initially, the Memory Bank is populated with the reference representation $\mathcal{I}_{r}$. We propose to condition the prediction for each target frame on the reference as well as on previous target predictions, by encoding in the Memory Bank the predicted representation $\mathcal{I}_{t}^{j}$, computed as in Eq. 1. This design transforms each intermediate frame into a pseudo-reference for subsequent frames. This objective discourages overfitting to individual image pairs and encourages robust, semantically grounded correspondences, forcing the model to disentangle semantics from low-level features. We supervise the predicted $\{\hat{y}^{j}_{t}\}_{j=1}^{J}$ with a Binary Cross-Entropy loss and a Dice loss [40].
@@ 3.4 Visualizing SANSA Feature Space
In this section we analyze how adaptation reshapes SAM2 feature space. In Fig. 4a, we extract features for both frozen SAM2 and SANSA for a set of objects belonging to unseen classes for our trained model, and visualize their first two Principal Components, labeled by class. Frozen SAM2 features exhibit weak class discriminability, reinforcing the hypothesis that semantic structure is entangled with task-specific features tailored for SAM2 original tracking objective, explaining the poor zero-shot performance of SAM2 in FSS. Consequently, the leading Principal Components reveal a mix of both semantic and non-semantic signals. Differently, our SANSA features show clear semantic discriminability, mapping novel classes into well-defined semantic clusters.
Figure 4: Semantic structure of feature space. (a) PCA visualization of frozen SAM2 and our SANSA features on a COCO fold with unseen classes, showing the first two principal components color-coded by class. SAM2 features exhibit weak semantic separability, indicating entanglement with other signals. (b) PCA-based RGB visualization of SANSA features across images with seen (e.g. umbrella) and unseen categories (e.g. person), showing consistent semantic mapping.
In Fig. 4b, we visualize SANSA object features via PCA, computed jointly across images, where the first three Principal Components are mapped to RGB channels. The visualization spans three images containing instances from both unseen (e.g., person, chair, ball) and seen (e.g., umbrella) categories. The consistent color mapping across instances of the same category highlights strong semantic grouping, demonstrating that similar concepts are mapped coherently in feature space despite visual variability.
We report an extended analysis in the Appendix.
@@ 4 Experiments
Implementation details. We employ SAM2 with Hiera-Large [52] as the feature encoder, incorporating AdaptFormer [11] in its last two layers. We note that SAM2 remains entirely frozen, with only the adapters (10M parameters) being trained. The model is optimized with Adam with a learning rate of 1e-4 for 5 epochs in strict few-shot segmentation and 20 epochs in the generalist setting.
Datasets. COCO-20i [43] is built on MSCOCO [34] and consists of 80 classes split into four folds, each with 20 classes. FSS-1000 [17] contains 1000 classes, with 520 for training, 240 for validation, and 240 for testing. LVIS-92i [37] is more challenging, selecting 920 classes from LVIS [18], divided in 10 folds. PASCAL-Part[37] includes four superclasses with 56 object parts across 15 classes. PACO-Part is built from PACO [47] and contains 303 classes, split in four folds.
@@ 4.1 Strict Few-Shot Segmentation Setting
We first evaluate our method in strict FSS setting. Following standard protocols [65, 56, 55], we train on base classes and evaluate on novel classes with k-shots. These experiment address the question of whether the semantic mapping learned on base classes can transfer meaningfully to novel categories.
@@
Table 1: Strict Few-Shot Segmentation setting. Results for $k$-shot segmentation on LVIS-92i, COCO-20i, and FSS-1000. We include both specialist models and approaches based on foundation models, trained and tested on disjoint classes. Zero-shot methods are also reported for reference.
@@
Method Venue Backbone Params few-shot segmentation LVIS-92i COCO-20i FSS-1000 1-shot 5-shot 1-shot 5-shot 1-shot 5-shot Zero-Shot Methods SAM 2 [48] - SAM2-L 224 M 16.5 26.3 32.2 44.2 73.0 84.3 Matcher [37] ICLR’24 DINOv2-L+SAM-H 945 M 33.0 40.0 52.7 60.7 87.0 89.6 GF-SAM [73] NIPS’24 DINOv2-L+SAM-H 945 M 35.2 44.2 58.7 66.8 88.0 88.9 Strict $k$-shot segmentation HSNet [41] ICCV’21 ResNet101 52 M 17.4 22.9 41.2 49.5 86.5 88.5 VAT[25] ECCV’22 ResNet101 46 M 18.5 22.7 41.3 47.9 90.3 90.8 AMFormer [65] NIPS’23 ResNet101 49 M - - 51.0 57.3 - - PerSAM [77] ICLR’24 SAM-H 641 M 11.5 - 23.0 - 71.2 - VRP-SAM [56] CVPR’24 RN50 + SAM-H 670 M 28.3 - 53.9 - 87.7 - SegIC [39] ECCV’24 DINOv2-G 1.2 B 40.5 - 53.6 - 88.5 - DiffewS [78] NIPS’24 StableDiffusion 890 M 33.9 43.7 52.2 60.7 90.2 90.6 SANSA this work SAM2-L 234 M 49.8 55.9 60.1 67.3 91.4 92.1
@@
Few-shot segmentation. In Tab. 1, we compare SANSA, besides specialist models, such as VAT [24] and HSNet [41], with the most relevant and recent approaches based on Foundation Models. These include methods that, like ours, leverage a single Foundation Model, such as SegIC [39] and DiffewS [78], as well as modular two-stage pipelines like VRP-SAM [56]. For completeness, we also report results from generalist zero-shot methods built on DINOv2 and SAM, including Matcher [37], GF-SAM [73], as well as our baseline, SAM2 in its zero-shot configuration.
@@
In the one-shot setting, SANSA consistently outperforms all prior methods, demonstrating superior generalization to unseen classes. Specifically, we surpass the best direct competitors SegIC, VRP-SAM and VAT by +9.3%, +6.2%, and +1.1% on LVIS-92i, COCO-20i, and FSS-1000, respectively. This performance gap remains consistent also against zero-shot approaches: compared to GF-SAM, SANSA achieves gains of +14.6%, +1.4%, and +3.4%. On the challenging LVIS-92i, including 920 fine-grained categories (e.g., bulldog, dalmatian), DINO-based methods (SegIC, GF-SAM, Matcher) tend to underperform, possibly due to DINO overly-semantic features [15, 76] (e.g., grouping distinct breeds under the concept of “dog”). Here, SANSA achieves a substantial +9.3% gain, suggesting that SAM2 encodes a latent hierarchical semantic structure, capturing both high-level semantic concepts (dominant in COCO-20i) and fine-grained distinctions (crucial for LVIS-92i).
@@
Regarding the 5-shot setting, we note that SegIC and VRP-SAM do not provide an inference pipeline to extend to $k$-shot. In contrast, SANSA natively models correspondences across multiple reference frames and it outperforms the best competitor, DiffewS, by +12.2% and +6.6% on LVIS-92i and COCO-20i, respectively. This performance gap remains consistent when compared to zero-shot models, with improvements of +11.7%, +3.2%, and +0.5% on LVIS-92i, FSS-1000, and COCO-20i.
@@
VRP-SAM SANSA(ours) Params 670M 234M point 38.4 53.4 +15.0 scribble 47.3 53.1 +5.8 box 49.7 54.3 +4.6 mask 53.9 60.1 +6.2
@@
Performance-Efficiency Trade-off. In Fig. 5, we analyze the trade-off between model size, inference speed (img/s), and performance. SANSA achieves state-of-the-art results while being the most lightweight solution. Specifically, it is: i) over 3× faster than the fastest competing method (GF-SAM), ii) more compact, introducing only adapter parameters on top of SAM2 (totaling 234M), and iii) substantially more accurate, outperforming GF-SAM by +9.3 mIoU on the challenging LVIS-92i benchmark. Importantly, SANSA keeps the SAM2 architecture entirely frozen, meaning that by storing only the adapter weights, it retains SAM2 state-of-the-art performance on Video Object Segmentation while also achieving top-tier results on Few-Shot Segmentation within a single model.
One-Shot Promptable Segmentation. By keeping SAM2 decoder frozen, our framework maintains its capability to generate masks from any prompt (i.e. points, scribbles, or boxes). As such, during inference, users can segment reference objects using simple points than requiring costly pixel-level masks. In LABEL:fig:promptable_combined we evaluate the performance with different prompts on the strict few-shot setting on COCO-20i against VRP-SAM [56], which also supports such prompts. SANSA outperforms VRP-SAM by +15.0%, +4.6%, and +5.8% when using points, scribbles, and boxes as annotations. More importantly, the performance drop from costly masks to point prompts is substantially smaller for SANSA (-6.7%) compared to VRP-SAM (-15.5%). We attribute this to the tightly coupled design of our architecture, which maximizes feature reuse by jointly leveraging the same representations for prompt encoding, feature matching, and mask decoding.
Table 2: Performance of SANSA against Generalist in-context models. Excluding zero-shot approaches, all methods are trained on COCO and ADE20k, and we report additional training datasets for each one. SegIC (*) uses additional supervision via a textual meta-prompt at test time.
few-shot segmentation part seg. Method Additional Datasets Params LVIS-92i COCO-20i FSS-1000 Pascal PACO 1-shot 5-shot 1-shot 5-shot 1-shot 5-shot Part Part Zero-Shot Methods (DINOv2 + SAM) Matcher n.a. 945M 33.0 40.0 52.7 60.7 87.0 89.6 42.9 34.7 GF-SAM n.a. 945M 35.2 44.2 58.7 66.8 88.0 88.9 44.5 36.3 Painter [63] NYUv2 354 M 10.5 10.9 33.1 32.6 61.7 62.3 30.4 14.1 SegGPT [64] Pascal, PACO 354 M 18.6 25.4 56.1 67.9 85.6 89.3 35.8 13.5 SINE [35] Object365, PACO 373 M 31.2 35.5 64.5 66.1 - - 36.2 - DiffewS [78] Pascal 890 M 31.4 35.4 71.3 72.2 87.8 88.0 34.0 22.8 SegIC [39] LVIS 1.2 B 47.8 - 74.5* - 88.4 - 31.2 18.7 SANSA - 234 M 36.2 42.5 76.4 77.3 88.2 89.0 40.4 29.2 SANSA LVIS 234 M 53.2 58.1 74.7 76.3 89.1 89.7 41.5 31.2 SANSA LVIS, PACO 234 M 51.8 57.6 74.1 75.5 89.4 89.9 49.6 42.8
@@ 4.2 Generalist In-Context Setting
Following recent few-shot segmentation works [39, 78, 37, 73], we evaluate SANSA in the in-context setting, where a single generalist model is tested across multiple benchmarks (cf. Tab. 2). Given the lack of a standardized training protocol, we explicitly report additional datasets used by each method beyond ADE20K and COCO, which are shared across all approaches. We evaluate three configurations of our approach: i) a minimal setup using only COCO and ADE20K, ii) an extended version incorporating LVIS, following the setup of SegIC [39], and iii) a variant including PACO to mitigate object-level bias and reinforce part segmentation capabilities.
When trained only on COCO and ADE20K coarse categories, SANSA exhibits strong out-of-domain generalization on the LVIS benchmark, outperforming DiffewS and SINE by +5.2% and +5.4%, respectively. Moreover, when LVIS is included in the training set (matching SegIC in-domain setup), SANSA further improves performance and surpasses SegIC by +5.4%. Despite not being trained at the part level, SANSA exhibits strong cross-task generalization capabilities, surpassing generalist baselines by a large margin (+7.5% and +8.4% over DiffewS, the best competitor on Pascal-Part and Paco-Part, respectively), including those trained on part segmentation datasets [64, 35]. Our method is only outperformed by zero-shot models, which, by design, are not biased by object-level training. To address this bias, we follow [64, 35] and augment our training set with PACO, leading to +6.5% on Paco-Part (in-domain) and +5.1% on Pascal-Part (out-of-domain) against the strong baseline of GF-SAM. Finally, we highlight that, within the generalist model category, SANSA is the most compact solution, with only 234M parameters.

### v2

@@ 3.2 SAM2 Feature Adaptation
We hypothesize that SAM2 features already encode semantic concepts, albeit entangled with signals specific to tracking, such as instance-level details and spatial biases. If such structure exists, then it should be learnable from a set of base classes by training few parameters [aghajanyan2020intrinsic, hu2022lora, rusu2018meta]. Thus, we opt for simple AdaptFormer [chen2022adaptformer] blocks, although our analysis is not tied to the adaptation method, as we will show in the Experimental Section. We integrate AdaptFormer blocks within the last two layers of the Image Encoder, as these encode higher-level semantic representations. Given down- and up-projection matrices $\mathbf{W}_{down}\in\mathcal{R}^{d,\tilde{d}},\mathbf{W}_{up}\in\mathcal{R}^{\tilde{d},d}$, an AdaptFormer block operates token-wise:
@@ 3.3 Training objective
Following standard practice in FSS, we adopt an episodic training paradigm [vinyals2016matching, shaban2017one, wang2019panet, lang2022learning]. We have access to a set of training episodes, each one containing annotated instances from a single class. We denote these episodes as $\{x^{i},y^{i}\}$, where $y^{i}\in[0,1]^{H\times W}$ is the mask that segments these instances.
@@
where $\{x_{r},a_{r}\}$ is the annotated reference image and $\{x^{j}_{t}\}_{j=1}^{J}$, are the $J$ unlabeled target images. We feed $\mathcal{M}_{train}$ to our model, which propagates the provided concept across target frames to predict the masklet $\{\hat{y}^{j}_{t}\}_{j=1}^{J}$. Initially, the Memory Bank is populated with the reference representation $\mathcal{I}_{r}$. We propose to condition the prediction for each target frame on the reference as well as on previous target predictions, by encoding in the Memory Bank the predicted representation $\mathcal{I}_{t}^{j}$, computed as in Eq.˜2. This design transforms each intermediate frame into a pseudo-reference for subsequent frames. This objective discourages overfitting to individual image pairs and encourages robust, semantically grounded correspondences, forcing the model to disentangle semantics from low-level features. We supervise the predicted $\{\hat{y}^{j}_{t}\}_{j=1}^{J}$ with a Binary Cross-Entropy loss and a Dice loss [milletari2016v].
@@ 3.4 Visualizing SANSA Feature Space
In this section we analyze how adaptation reshapes SAM2 feature space. In Fig.˜4a, we extract features for SAM2 and SANSA for a set of objects belonging to unseen classes for our trained model, and visualize their first two Principal Components (PCs), labeled by class. Frozen SAM2 features show weak class discriminability, reinforcing the hypothesis that semantic structure is entangled with task-specific features tailored for the original tracking objective, explaining SAM2 poor FSS performance. Consequently, the leading PCs reveal a mix of semantic and non-semantic signals. In contrast, our SANSA features show clear semantic discriminability, mapping novel classes into well-defined semantic clusters.
Figure 4: Semantic structure of feature space. (a) PCA visualization of frozen SAM2 and our SANSA features on a COCO fold with unseen classes, showing the first two principal components color-coded by class. SAM2 features exhibit weak semantic separability, indicating entanglement with other signals. (b) PCA-based RGB visualization of SANSA features across images with seen and unseen categories, showing consistent semantic mapping. (c) Part-level semantics and cross-image consistency. We cluster features per object and match clusters across image pairs via Hungarian Matching. This reveals that SANSA captures fine-grained distinctions (e.g., handlebar vs. wheel), spatial layout (e.g., upper vs. lower wheel), and produces representations that align across images.
In Fig.˜4b, we visualize SANSA features using PCA, computed jointly across images, where the first three PCs are mapped to RGB channels. The visualization spans three images containing instances from both unseen (e.g., person, chair, ball) and seen (e.g., umbrella) categories. The consistent color mapping across instances of the same category highlights strong semantic grouping, mapping similar concepts coherently in feature space despite visual variability.
Finally, in Fig.˜4c, we shift from coarse category level to fine-grained correspondences at the part level. Following [zhang2023tale], we cluster object features with K-Means and match centroids across image pairs via the Hungarian Algorithm, then visualize matched clusters with shared colors to assess semantic consistency. The results show that, despite not being trained with part-level supervision, SANSA features encode part-level understanding (e.g. hand vs arm, handlebar vs wheel), spatial layout (e.g. upper vs lower wheel), and produce representations that align across images.
We provide an in-depth study of semantic representations in Appendix˜B, using PCA, clustering, and linear probing to analyze how semantics are encoded in SAM2 and made explicit through adaptation.
@@ 4 Experiments
Implementation details. We employ SAM2 with Hiera-Large [ryali2023hiera] as encoder. AdaptFormer [chen2022adaptformer] is inserted into the last two blocks, with hidden size set to $0.3\times$ the block channel dimension in the strict few-shot setting and $0.8\times$ in the generalist. SAM2 is frozen, and only the adapters are trained ($\sim$10M params in the strict case and $\sim$25M in the generalist). We train with AdamW and learning rate $10^{-4}$ for 5 epochs (strict) and 20 (generalist), with k=1k{=}1 (a single annotated reference) and sequence length J=3J{=}3. The same model is evaluated on 1-shot and 5-shot. Full details are in Appendix˜H.
Datasets. COCO-20i [nguyen2019feature] is built on MSCOCO [lin2014microsoft] and consists of 80 classes split into four folds, each with 20 classes. FSS-1000 [gao2022mutually] contains 1000 classes, with 520 for training, 240 for validation, and 240 for testing. LVIS-92i [liu2023matcher] is more challenging, selecting 920 classes from LVIS [gupta2019lvis], divided in 10 folds. PASCAL-Part [liu2023matcher] includes four superclasses with 56 object parts across 15 classes. PACO-Part is built from PACO [ramanathan2023paco] and contains 303 classes, split in four folds.
@@ 4.1 Strict Few-Shot Segmentation Setting
We first evaluate our method in strict FSS setting. Following standard protocols [wang2023amf, sun2024vrp, shi2022dense], we train on base classes and evaluate on novel classes with k-shots. These experiment address the question of whether the semantic mapping learned on base classes can transfer meaningfully to novel categories.
@@
Table 1: Strict Few-Shot Segmentation setting. Results for $k$-shot segmentation on LVIS-92i, COCO-20i, and FSS-1000. We include both specialist models and approaches based on foundation models, trained and tested on disjoint classes. Training-free methods are also reported for reference.
@@
Method Venue Backbone Params few-shot segmentation LVIS-92i COCO-20i FSS-1000 1-shot 5-shot 1-shot 5-shot 1-shot 5-shot Training-free methods SAM 2 [ravi2024sam] - SAM2-L 224 M 16.5 26.3 32.2 44.2 73.0 84.3 PerSAM [zhang2023personalize] ICLR’24 SAM-H 641 M 11.5 - 23.0 - 71.2 - Matcher [liu2023matcher] ICLR’24 DINOv2-L+SAM-H 945 M 33.0 40.0 52.7 60.7 87.0 89.6 GF-SAM [zhang2024bridge] NeurIPS’24 DINOv2-L+SAM-H 945 M 35.2 44.2 58.7 66.8 88.0 88.9 Strict $k$-shot segmentation AMFormer [wang2023amf] NeurIPS’23 ResNet101 49 M - - 51.0 57.3 - - HMNet [xu2025hybrid] NeurIPS’24 RN50 39 M - - 52.1 58.9 - - VRP-SAM [sun2024vrp] CVPR’24 RN50 + SAM-H 670 M 28.3 - 53.9 - 87.7 - SegIC [meng2024segic] ECCV’24 DINOv2-G 1.2 B 40.5 - 53.6 - 88.5 - DiffewS [zhu2024unleashing] NeurIPS’24 StableDiffusion 890 M 33.9 43.7 52.2 60.7 90.2 90.6 SANSA NeurIPS’25 SAM2-L 234 M 48.8 53.9 60.2 64.3 91.4 92.1
@@
Few-shot segmentation. In Tab.˜1, we compare SANSA, besides specialist models, such as AMFormer [wang2023amf] and HMNet [xu2025hybrid], with the most relevant and recent approaches based on Foundation Models. These include methods that, like ours, leverage a single Foundation Model, such as SegIC [meng2024segic] and DiffewS [zhu2024unleashing], as well as modular two-stage pipelines like VRP-SAM [sun2024vrp]. For completeness, we also report results from generalist training-free methods built on DINOv2 and SAM, including PerSAM [zhang2023personalize], Matcher [liu2023matcher], GF-SAM [zhang2024bridge], as well as our baseline, i.e. frozen SAM2.
@@
In the one-shot setting, SANSA consistently outperforms all prior methods, demonstrating superior generalization to unseen classes. Specifically, we surpass the best direct competitors SegIC, VRP-SAM and DiffwS by +8.3%, +6.3%, and +1.2% on LVIS-92i, COCO-20i, and FSS-1000, respectively. This performance gap remains consistent also against training-free approaches: compared to GF-SAM, SANSA achieves gains of +13.6%, +1.5%, and +3.4%. On the challenging LVIS-92i, including 920 fine-grained categories (e.g., bulldog, dalmatian), DINO-based methods (SegIC, GF-SAM, Matcher) tend to underperform, possibly due to DINO overly-semantic features [engstler2023understanding, zhang2023tale] (e.g., grouping distinct breeds under the concept of “dog”). Here, SANSA achieves a substantial +8.3% gain, suggesting that SAM2 encodes a latent hierarchical semantic structure, capturing both high-level semantic concepts (dominant in COCO-20i) and fine-grained distinctions (crucial for LVIS-92i).
@@
Regarding the 5-shot setting, we note that SegIC and VRP-SAM do not provide an inference pipeline to extend to $k$-shot. In contrast, SANSA natively models correspondences across multiple reference frames and it outperforms the best competitor, DiffewS, by +10.2% and +3.6% on LVIS-92i and COCO-20i, respectively. Compared to training-free models, SANSA shows improvements of +9.7%, +3.2%, on LVIS-92i and FSS-1000, while suffering a -2.5% gap on COCO-20i w.r.t. GF-SAM. We also report results of upgrading SAM-based baselines to SAM2 in Appendix˜D.
@@
VRP-SAM SANSA(ours) Params 670M 234M point 38.4 53.4 +15.0 scribble 47.3 53.1 +5.8 box 49.7 54.3 +4.6 mask 53.9 60.2 +6.3
@@
Performance-Efficiency Trade-off. In Fig.˜5, we analyze the trade-off between model size, inference speed (img/s), and performance. SANSA achieves state-of-the-art results while being the most lightweight solution. Specifically, it is: i) over 3× faster than the direct competing method (GF-SAM), ii) more compact, introducing only adapter parameters on top of SAM2 (totaling 234M), and iii) substantially more accurate, outperforming GF-SAM by +13.6% mIoU on the challenging LVIS-92i benchmark. Importantly, SANSA keeps the SAM2 architecture entirely frozen, meaning that by storing only the adapter weights, it retains SAM2 state-of-the-art performance on Video Object Segmentation while also achieving top-tier results on few-shot segmentation within a single model. A large-scale annotation study quantifying speed/quality trade-offs is presented in Appendix˜C.
One-Shot Promptable Segmentation. Our framework keeps SAM2 decoder frozen, maintaining its capability to generate masks from any prompt (i.e. points, scribbles, or boxes). As such, during inference, users can segment reference objects by providing a simple point, without requiring costly pixel-level masks. In Fig.˜6 we evaluate the performance with different prompts on strict FSS setting on COCO-20i against VRP-SAM, which also supports such prompts. SANSA shows gains of +15.0%, +4.6%, and +5.8% when using points, scribbles, and boxes as annotations. More importantly, the performance drop from masks to point prompts is substantially smaller for SANSA (-6.8%) compared to VRP-SAM (-15.5%). We attribute this to the tightly coupled design of our architecture, which maximizes feature reuse by jointly leveraging the same representations for prompt encoding, feature matching, and mask decoding. Prompt generation details and qualitative results are in Appendix˜E.
@@ 4.2 Generalist In-Context Setting
Following recent few-shot segmentation works [meng2024segic, zhu2024unleashing, liu2023matcher, zhang2024bridge], we evaluate SANSA in the in-context segmentation setting, where a single generalist model is tested across multiple benchmarks (cf. Tab.˜2). Given the lack of a standardized training protocol, we explicitly report additional datasets used by each method beyond ADE20K and COCO, which are shared across all approaches. We evaluate three configurations of our approach: i) a minimal setup using only COCO and ADE20K, ii) an extended version incorporating LVIS, following the setup of SegIC [meng2024segic], and iii) a variant including PACO to mitigate object-level bias and reinforce part segmentation capabilities.
When trained only on COCO and ADE20K coarse categories, SANSA exhibits strong out-of-domain generalization on the LVIS benchmark, outperforming DiffewS and SINE by +4.8% and +5.0%, respectively. Moreover, when LVIS is included in the training set (matching SegIC in-domain setup), SANSA further improves performance and surpasses SegIC by +5.4%. Despite not being trained at the part level, SANSA exhibits strong cross-task generalization capabilities, surpassing generalist baselines by a large margin (+7.5% and +8.4% over DiffewS, the best competitor on Pascal-Part and Paco-Part, respectively), including those trained on part segmentation datasets [wang2023seggpt, liu2024simple]. Our method is only outperformed by training-free models, which, by design, are not biased by object-level training. To address this bias, we follow [wang2023seggpt, liu2024simple] and augment our training set with PACO, leading to +6.7% on Paco-Part (in-domain) and +4.6% on Pascal-Part (out-of-domain) against the strong baseline of GF-SAM. Finally, we highlight that, within the generalist model category, SANSA is the most compact solution, with only 250 M parameters.
Table 2: Performance of SANSA against Generalist In-context models. Excluding training-free approaches, all methods are trained on COCO and ADE20k, and we report additional training datasets for each one. SegIC (*) uses additional supervision via a textual meta-prompt at test time.
few-shot segmentation part seg. Method Additional Datasets Params LVIS-92i COCO-20i FSS-1000 Pascal PACO 1-shot 5-shot 1-shot 5-shot 1-shot 5-shot Part Part Training-free methods Matcher [liu2023matcher] n.a. 945M 33.0 40.0 52.7 60.7 87.0 89.6 42.9 34.7 GF-SAM [zhang2024bridge] n.a. 945M 35.2 44.2 58.7 66.8 88.0 88.9 44.5 36.3 Painter [wang2023images] NYUv2 354 M 10.5 10.9 33.1 32.6 61.7 62.3 30.4 14.1 SegGPT [wang2023seggpt] Pascal, PACO 354 M 18.6 25.4 56.1 67.9 85.6 89.3 35.8 13.5 SINE [liu2024simple] Object365, PACO 373 M 31.2 35.5 64.5 66.1 - - 36.2 - DiffewS [zhu2024unleashing] Pascal 890 M 31.4 35.4 71.3 72.2 87.8 88.0 34.0 22.8 SegIC [meng2024segic] LVIS 1.2 B 47.8 - 74.5* - 88.4 - 31.2 18.7 SANSA - 250 M 36.2 42.5 76.4 77.1 88.2 89.0 40.4 29.2 SANSA LVIS 250 M 53.2 58.1 74.7 76.3 89.1 89.7 41.5 31.2 SANSA LVIS, PACO 250 M 50.3 59.0 75.6 78.6 90.0 91.0 49.1 43.0
4.3 Generalization across domains and styles
To explore how SANSA generalizes beyond the scope of standard benchmarks, we follow [sun2024vrp, liu2023matcher] and present qualitative examples drawn from in-the-wild image pairs, shown in Fig.˜7. In each case, the model is tasked with few-shot segmentation using a reference image from COCO [lin2014microsoft] and a target image collected from the web, offering a complementary view to benchmark evaluations.
These examples introduce a different challenge compared to traditional benchmarks, focusing on domain and style shifts, such as when the target image is a cartoon or stylized sketch, or when the visual appearance changes dramatically between the reference and target. These results demonstrate SANSA strong generalization. Importantly, this robustness across domains and styles is inherited from the frozen SAM2. Our method preserves this capability by learning a semantic mapping within the frozen feature space, enabling reliable correspondence even under significant distribution shifts.